Redesigned Computers has 7 percent (APR) coupon bonds outstanding with a current market price of $882.80. The yield to maturity is 8.34 percent (APR) and the face value is $1,000. Interest is paid semiannually. How many years is it until this bond matures?

Please show how to use the financial calculator to solve it.
Answer #1

Information provided:

Face value= future value= $1,000

Coupon rate= 7%/2= 3.5%

Coupon payment= 0.035*1,000= $35

Yield to maturity= 8.34%/2= 4.17% per semi-annual period

Current price= present value= $882.80

The years to maturity is computed by entering the below in a financial calculator:

FV= 1,000

PMT= 35

I/Y= 4.17

PV= -882.80

Press the CPT key and N to compute the years to maturity.

The value obtained is 31.9987.

Therefore, the years to maturity is 31.9987/2= 15.9991 years \rightarrow 16 years.  

In case of any query, kindly comment on the solution.